OpenStudy (anonymous):

Find the slope m of the line that passes through the points (3, –8) and (–7, –4). (Points : 5) slope m = –5 /2 slope m = –2 /5 slope m = 2 /5 slope m = 5 /2

OpenStudy (anonymous):

m = rise/ run = (-4-(-8))/(-7-3) = 4/10 = 2/5 There! =D

OpenStudy (anonymous):

can you show me how to do it

OpenStudy (anonymous):

yup sure. The slope is rise/run so you must find the y-value of the first point - the y value of the second point. Then divide that by the x-value of the first point - x value of second point. That way, you can see how much the y value changes when the x value changes by 1 which is the slope. Tell me if it doesn't make sense.